FTIR ISOTOPIC AND DFT STUDIES OF TRANSITION METAL-CARBON CLUSTERS CONDENSED IN SOLID ARGON: CrC_{3}

\underline{SARAH A. BATES}, C. M. L. RITTBY, and W. R. M. GRAHAM, Department of Physics and Astronomy, Texas Christian University, Fort Worth, TX 76129.

We report here the first results from Fourier transform infrared (FTIR) and density functional theory (DFT) studies that have been initiated on the structures and vibrational fundamentals of transition metal-carbon clusters, including chromium-carbon (Cr_{n}C_{m}) species. CrC_{3} has been produced by ablating carbon and chromium rods with a Nd:YAG laser and trapping the products in solid Ar at ∼10 K. Earlier gas phase work using photoelectron spectroscopy and DFT calculations had given evidence of both C_{2v} (fan-shaped) and linear isomers for CrC_{3}.^{a} In the present work, extensive FTIR measurements of vibrational frequencies and ^{13}C isotopic shifts measurements compared with the predictions of DFT calculations at the B3LYP/6-311G+(3df ) level confirm the assignment of the ν_{1}(σ) fundamental of the linear isomer of CrC_{3} at 1789.5 cm^{−1}.

aH. -J. Zhai, L. -S. Wang, P. Jena, G. L. Gustev, and C. W. Bauschlicher, Jr., J. Chem. Phys. \underline{120}, 8996 (2004).